Convert the following:show your work(1 point each)        

  1. 10-kilograms equals _____ pounds
  2. 12.5 ml is equal to _____ teaspoons
  3. 45 ml is equal to _____ Tablespoons
  4. 4 tsp is equal to _____ ml
  5. 3 Tbsp is equal to _______ounces
  6.   0.1 grams is equal to _______ mcg
  7. 1 cup is equal to ____ounces
  8. 1 liter is equal to _____ ml

  1. 3 Tbsp is equal to ___ tsp

  1. 75 lbs.is equal to _____kg.

  1. 9 oz is equal to _______ml

  1. 38 degrees Celsius is equal to_________ Fahrenheit.

  1. 97 degrees Fahrenheit is equal to________ Celsius.

Round the following numbers as directed (1 point each)

            500/34 (round to the nearest hundredth) _____________

            350/21 (round to the nearest tenth) _____________

            403/10 (round to the nearest whole number) ______________

Matching or multiple choice (1 point each)

  1. Match the following scenarios to the rights they uphold:
    ____Right medication

____Right amount

____Right time

____Right route

____Right patient

____Right documentation

____Right to refuse

____Right room

  1. Not one of the seven rights
  2. Checking the patient’s identification band and asking the client to state his name
  3. Consistently giving a daily “ac” medication before the client eats breakfast
  4. Carefully compare the dosage ordered with the dose on hand (the unit-dose packet)
  5. Holding a medication if a patient is nauseated and doesn’t want to take it
  6. Calling to question giving medication orally when the client has trouble swallowing
  7. Comparing the drug ordered with the label on the package, bottle, or unit-dose pack
  8. Recording the time and dose of an IM pain medication with the route, and location

Solve the following dosage calculations by using any of the 3 methods. **Show your work and label your answer. Make sure you round appropriately. 1 pt each    

The order is for 60mg. The label reads gr. 1 per tablet. You would give _______tablet(s)

The order is for 200 mg. The label reads 250mg /5ml. You would give _________ml(s).

The order is for 60 ml. You would give ________tablespoon(s).

The order is for 60mg. The label reads 60 mg/2 ml. You would give _______ml.

           

The order is for 1000 ml. of I.V. solution to be infused at 125 ml/hr. You would infuse the solution for ________hours.

           

The order is for 6000 ml of IV solution to run over a 24-hour period. The drop factor of the I.V tubing is 10 gtts/ml. You would infuse _________gtts/min.

           

Your patient weighs 75 kg and you are ordered to infuse 250 mg dobutamine in 500 ml NS at 10 mcg/kg/min. How many milligrams of dobutamine will infuse per hour? _______

           

The order is Solumedrol 3mg/kg for a child weighing 20kg. Solumedrol is available as 125 mg / 2 ml. How many ml must the nurse administer?__________

The order:KayCiel 30 mEq PO Bid pc. Available: Kay Ciel 45 mEq/15 ml. How many ml of KayCiel should you administer? ____________

           

The order is for 25 mg. The label reads 75mg/ml. How many ml(s) would you give? _______

           

The order is for 1.5 g the label reads 1 tablet equals 3 g. How many tablet (s) would you give? __________

The order is for 400mcg of thyroxin. The label reads 0.2 mg tablets. How many tablets would you give? ____________

Mrs. N is taking Motrin 600mg PO q6hr around the clock for joint pain. How many grams of Motrinis Mrs. N consuming daily? ___________

The order is Penicillin G 1.2 million units IM daily. Available: Penicillin G 9 million units/1 ml. How many ml will you give? __________

Mrs. J has a history of pernicious anemia and visits her physician’s office bi-monthly for vitamin B12 injections. Order: vitamin B12 0.50 mg IM bi-monthly. Supply: vitamin B12 1000mcg/mL. How many mL of vitamin B12 will Mrs. J be receiving with each dose? __________

Mrs. A is admitted to the outpatient care center for a blood transfusion for a Hgb of 7.2. Order: PRBC 1 unit (480 ml) to infuse over 4 hours via an infusion pump. What is the flow rate in mL/h? __________

After reviewing Mr. B's morning labs, his physician orders an SQ dose of Heparin to be given STAT. Order: Heparin 6,000 units SQ STAT. Supply: Heparin 10,000 units/mL. How many mL of Heparin will Mr. B be given? _____________

Mr. R is a 74 y.o.man being seen in the Acute Care Community Clinic. He is diagnosed with pneumonia and the physician orders a STAT dosage of Cefepime to be given in the clinic before he is admitted to the hospital. Order: Cefepime 750 mg IM q12h - first dose now. Supply: Cefepime 300 mg/mL. How many mL will Mr. R be receiving intramuscularly with each dose? ___________

Ms. W is not able to tolerate oral fluids secondary to food poisoning. Her IV fluids are ordered to infuse at 150 mL/h via infusion pump. What will you expect her 12-hour IV fluid intake to be in mL? ___________

Mr. Q is a 56 y.o.male admitted to the hospital after a motor vehicle accident. His admission weight is 280 lbs. What is his admission weight in kg? ______________

Ms. F is being converted from an IV antibiotic to PO Cefadroxil in preparation for her discharge home. Order: Cefadroxil 0.5 g PO bid. Supply: cefadroxil 500 mg tablets. How many tablet(s) will Ms. F be taking in a single dose? ______________

The client has an order for lorazepam (Ativan) 0.5 mg IV q6h PRN for agitation. Available is a vial containing 2 mg/ml. The nurse should draw up ____ ml of solution for injection. _______________

A client has a temperature of 101.2 F. There is an order for acetaminophen (Tylenol) 650 mg PO for fever, and 325 mg tabs are available. The nurse should give ______ tablets.
Answer #1

Convert the following

a.10kg=..pounds

1kg=2.205 pounds

So 10 kg =???

2.205×10

=22 pounds

b.12.5ml= ..teaspoon

1 ml=0.203 US teaspoon so 12.5 ml=??

12.5×0.203=2.5 teaspoon

c.1ml=0.0676 US tablespoon

So 45ml =??

45×0.0676=3 tablespoon

d.1 teaspoon =4.929ml

4×4.929=19.7 ml

e.1 tablespoon =0.5 ounce

Then 3 tablespoon=??

3×0.5=1.5 ounce

f.1 gram=1000000mcg

So 0.1 gram=??

=0.1×1000000÷1

=100000mcg

g.1 cup=8 ounce (1×8)

h.1 L=1000ml

i.1 tablespoon =3 teaspoon

So 3 tablespoon=??

3×3=9 teaspoon

j.1 pound=0.454kg

75lbs=??

0.454×75=34kg

k.1 ounce=29.574

9oz=??

29.574×9=266ml

ax.F=(0°C×9/5)+32

38=(C×9/5)+32

=100.4°F

all.C=F_32×5÷7

=F=97°

=97_32×5÷7

=65×5÷7

=36°C

Rounding off

●500/34=14.70

=15

●350/21

=16.66

=17

●403/10

=40.3

=40